### v1

# On the robustness of modeling grounded word learning through a child’s egocentric input

Abstract
What insights can machine learning bring to understanding human language acquisition? Large language and multimodal models have achieved remarkable capabilities, but their reliance on massive training datasets creates a fundamental mismatch with children, who succeed in acquiring language from comparatively limited input. To help bridge this gap, researchers have increasingly trained neural networks using data similar in quantity and quality to children’s input. Taking this approach to the limit, Vong et al., (2024) showed that a multimodal neural network trained on 61 hours of visual and linguistic input extracted from just one child’s developmental experience could acquire word-referent mappings. However, whether this approach’s success reflects the idiosyncrasies of a single child’s experience, or whether it would show consistent and robust learning patterns across multiple children’s experiences was not explored. In this article, we applied automated speech transcription methods to the entirety of the SAYCam dataset, consisting of over 500 hours of video data spread across all three children. Using these automated transcriptions, we generated multi-modal vision-and-language datasets for both training and evaluation, and explored a range of neural network configurations to examine the robustness of simulated word learning. Our findings demonstrate that networks trained on automatically transcribed data from each child can acquire and generalize word-referent mappings across multiple network architectures. These results validate the robustness of multimodal neural networks for grounded word learning, while highlighting the individual differences that emerge in how models learn when trained on each child’s developmental experiences.
1 Introduction
Recent progress in large language models and multi-modal AI systems have led to fluent language generation and accurate object recognition, among other human-like abilities. These developments have naturally piqued the interest of many cognitive scientists and developmental psychologists, raising the question of whether these computational advances can help explain how children acquire language in the first few years of their lives. Yet, there is a striking asymmetry to reckon with: can machines trained to learn language from billions, even trillions, of word tokens provide insight into how children learn language from mere millions (Frank,, 2023)? Addressing this disparity requires examining whether these language learning mechanisms can succeed under more realistic developmental constraints, combining these computational methods with richer and more ecologically valid datasets from developmental psychology.
Recent efforts towards this goal have examined what can be learned from linguistic (and other sensory) data that is comparable in quantity or kind to what children receive naturally (Huebner et al.,, 2021; Wang et al.,, 2023; Warstadt et al.,, 2024; Qin et al.,, 2023; Vong et al.,, 2024; Hu et al.,, 2024; Zhuang et al.,, 2024). For example, Huebner et al., (2021) examined language learning in neural networks from aggregated child-directed speech corpora (MacWhinney,, 1992), demonstrating that models can acquire syntactic and semantic knowledge in the limited-data regime that is more representative of the kinds of input children receive. Similarly, the BabyLM challenge (Warstadt et al.,, 2024) examined what kinds of linguistic competencies can be learned from training language models on datasets consisting of 10M to 100M words (Hu et al.,, 2024; Charpentier et al.,, 2025). Separately, Zhuang et al., (2024) studied how visual information can improve language learning, especially when using developmentally plausible amounts of data.
Additional recent work has aimed to increase the realism of the training data by using egocentric, head-mounted camera recordings from developing children, such as SAYCam (Sullivan et al.,, 2022) and BabyView (Long et al.,, 2024). These datasets provide a unique glimpse into what children actually see and hear over the course of language learning, providing a testbed for what can be learned by training models from this kind of naturalistic, high-quality input. Researchers have shown that certain kinds of syntax and semantics, as well as word-referent mappings, are all learnable, even under the extreme constraint of only using data collected from a single child (Wang et al.,, 2023; Qin et al.,, 2023; Vong et al.,, 2024; Long et al.,, 2024).
In recent work most related to the current article, Vong et al., (2024) demonstrated that a multimodal neural network trained on vision-and-language data extracted from a single child could acquire word-referent mappings. Specifically, they used a subset of a single child in SAYCam Sullivan et al., (2022), curating a dataset of paired child-directed utterances with corresponding video frames from 61 hours of video. This dataset was used to train a multimodal neural network utilizing two separate encoders for visual and linguistic inputs, combined with a contrastive objective to align embeddings from both modalities into a shared multimodal representation space. The model’s knowledge of word-referent mappings was evaluated on two tasks involving 22 within-distribution categories and 64 out-of-distribution categories, achieving high performance on the former and modest performance on the latter. This work demonstrated that grounded word-referent mappings could be acquired from a subset of one child’s experience from relatively generic neural networks. However, a major limitation is that its findings are reliant on the data from a single child (as only one child’s data had transcripts at the time), raising questions about how robust or generalizable the findings are.
In this article, we examine the robustness of the findings reported in Vong et al., (2024) by utilizing the full 500 hours of video from the SAYCam dataset (Sullivan et al.,, 2022) and training distinct models on each child’s data. The main contributions of our study are:
  • We use automated speech recognition methods to fully annotate and transcribe approximately 500 hours of video from the SAYCam dataset, leading to multimodal datasets consisting of over 1.77M video frames paired with 266K utterances, a dataset 7 times larger than the previously available manually transcribed dataset.
  • We assess the robustness of learning word-referent mappings from naturalistic visual-linguistic data by training separate vision-language models on each of the three different children in the SAYCam dataset from automated transcripts. Our results demonstrate how word-referent mappings are learnable across different datasets, despite varying degrees of visual-linguistic alignment from child to child.
  • We also explore robustness across different language model configurations and training objectives, finding modest variations in performance, suggesting that dataset characteristics outweigh specific modeling or architectural choices.
2 Datasets
Our approach requires datasets that closely mirror the visual and linguistic experience from the perspective of developing infants. One dataset that fits this criterion is SAYCam(Sullivan et al.,, 2022), a video dataset capturing the egocentric, longitudinal perspective of 3 different children aged 6 to 32 months, collected via head-mounted cameras. While previous attempts to train language and multimodal models from SAYCam showed promise in answering questions around learnability (Wang et al.,, 2023; Vong et al.,, 2024; Qin et al.,, 2023), their results were limited by the availability of manually transcribed natural language data, primarily to a subset of one of the three children’s data. In this work, we expand the amount of available data for training by a factor of 7 through the use of automated speech transcription tools (Radford et al.,, 2023; Bain et al.,, 2023). This enabled the creation of new training and evaluation datasets for each of the three children in SAYCam, providing an opportunity to replicate our studies of simulated word learning with multiple children’s data.
2.1 Training Datasets
Transcribing videos. In order to train separate vision-language models on each child’s data, we needed a way to transcribe all of the speech from the raw videos. To efficiently transcribe the long-form audio from 500 hours of video from SAYCam, we used WhisperX (Bain et al.,, 2023), a tool which leverages the transcription capabilities of OpenAI’s Whisper model (Radford et al.,, 2023) (using the v3-large model with the language set to English), and applies further processing for efficient transcription and speaker diarization and word-level timestamps. This provided us with the transcribed text of each separate utterance, the start and end timestamps, and the predicted speaker label. Additional details surrounding the automatic transcription procedure and evaluation for speech transcription accuracy can be found in Zhang et al., (2025).
Due to the inherent challenges of recording audio from a child’s head-mounted camera, the audio quality, especially for child-directed speech is highly variable across videos. Nevertheless, an analysis comparing a subset of manually vs. automatically transcribed speech content showed that it was relatively accurate for adult speech and also sufficiently similar to the previously manually transcribed subset (Zhang et al.,, 2025). Furthermore, although the relative timings for the detected start and end timestamps also showed some discrepancies, this is not necessarily a major issue: Vong et al., (2024) observed that imperfect timing information (in the manual transcriptions) were still sufficient for models to acquire word-referent mappings.
Whisper was not a panacea for annotating the data: it showed relatively poor performance on diarization accuracy (correctly distinguishing between different speakers in a video). Therefore, we did not attempt to distinguish between different speakers (e.g. child-directed versus child-produced speech), and retained all of the transcribed utterances. Another limitation was from hallucinations in the Whisper transcriptions, which resulted in instances of repetitive utterances, [^1] This included filtering repetitive phrases within a single utterance, as well as filtering adjacent utterances with the same content. In both cases, we only kept a single instance of the repeated utterance., requiring them to be filtered out. Additional pre-processing steps included removing all punctuation and lower-casing the transcribed text. Tokenization was performed at the word level using spaCy (Honnibal and Montani,, 2017) and was subsequently used to create separate tokenizers from the transcribed vocabulary of each dataset split (after excluding any words with a frequency of two or fewer). After filtering and pre-processing, this resulted in 266K automatically transcribed utterances across the entire SAYCam dataset.
Training datasets. To create vision-language datasets for training, we used this set of transcribed utterances to extract video frames in their original resolution that were temporally aligned with each utterance. For each detected utterance, we used the detected starting timestamp to extract up to 16 frames from the video at a rate of 3.75 fps, corresponding to 4.27s of video. Each set of extracted of video frames was paired with their corresponding utterance, forming the basis of our vision-language dataset for training. The total amount of available training data from automating the transcription process is roughly 7 times larger than what was previously available for a single baby via manual transcription (Sullivan et al.,, 2022; Vong et al.,, 2024). We split the utterances into five different dataset splits, making sure that each split uses data from at most one child. The five different splits are:
  • S-Whisper-2022: This split is based on the subset of the Whisper utterances from the same set of videos that were previously manually transcribed for baby S, as described in Sullivan et al., (2022).
  • S-Whisper-Disjoint: This split uses the complement of videos from S-Whisper-2022, the set of transcribed utterances from new videos from baby S that were not previously manually transcribed.
  • S-Whisper: This split uses the full set of Whisper utterances from baby S’s videos.
  • A-Whisper: This split uses the full set of Whisper utterances from baby A’s videos.
  • Y-Whisper: This split uses the full set of Whisper utterances from baby Y’s videos.
For each dataset split, videos were partitioned into independent train, validation, and test sets. Summary statistics for each of these splits are shown in Table 1.
2.2 Evaluation Datasets
Figure 1: Evaluation frames from Labeled-S-V2, Labeled-A and Labeled-Y. Here we present three randomly selected evaluation frames from four different evaluation categories (ball, toy book and cereal), with each row indicating frames derived from each of the three children in the SAYCam dataset.
For model evaluation, we primarily focus on image classification tasks, which are based on the looking-while-listening paradigm from early word learning studies (Golinkoff et al.,, 1992). In this evaluation procedure, models are presented with a target word along with four video frames (one belonging to the target category, and three distractors from other randomly sampled categories). The task is to select the video frame corresponding to the target word, where these labeled frames are sourced from SAYCam directly. We test our models on a number of different evaluation datasets:
  • Labeled-S: This first evaluation set consists of 22 different visual categories derived from baby S’s egocentric video data, covering the most common kinds of objects and places both seen and spoken about (Orhan et al.,, 2020; Vong et al.,, 2024). For each category, there are 100 separate evaluation trials for a total of 2200 evaluation trials. The full set of categories can be found in Appendix C.
  • Labeled-S-V2, Labeled-A and Labeled-Y: In addition to Labeled-S, we create additional evaluation datasets based on the datasets from each child, using a mixture of automatic and manual filtering (see Appendix A for additional details). For each category, there are 100 separate evaluation trials, for a total of 4100 evaluation trials. The frames used for evaluation were (primarily) derived from the test splits from each baby. [^2] For Labeled-S-V2, due to some differences in dataset splitting, there was some partial overlap between the videos used for extracting frames for training and evaluation. These 41 categories cover common visual concepts encountered by each child in SAYCam, including body parts (arm, hair, knee), clothing (shirt, jeans, shoe), living spaces (kitchen, porch, table), objects (ball, bottle, boy) and food (banana, cereal). A limited set of evaluation frames are shown in Figure 1, and the full set of categories can be found in Appendix C.
3 Models
Figure 2: Model architectures. We explore three different multi-modal architectures. The first two architectures utilize a contrastive loss with either a simple Embedding layer with averaging across words (CVCL), or a 2-layer Transformer Decoder (CVCL+T). The third is also a 2-layer Transformer Decoder, but incorporates an additional language modeling loss head (CVCL+T+LM), with an additional weight parameter to balance the two losses. All models use a pre-trained Vision Transformer as their vision encoder, trained only from the visual data from each child, along with a learned, absolute positional embedding scheme in the language encoder.
We begin by describing the CVCL (Child’s View for Contrastive Learning) model from Vong et al., (2024) before describing the variants we explore in this work. As described in the previous section, the input provided to the model is in the form of pairs of video frames and child-directed utterances. From this input, we need a method to align the representations from image frames with the representations of corresponding utterances, thereby linking visual referents to their corresponding words. To accomplish this feat, CVCL trains two neural networks jointly, learning to match image frames with their corresponding utterances. One of these networks (the image encoder) embeds video frames, while the other network (the language encoder), separately embeds utterances, projecting both into a shared multimodal representation space. This operationalizes the cross-situational learning approach from cognitive science Yu and Smith, (2007) through the lens of modern neural networks.
Specifically, CVCL’s vision encoder is designed to take in static video frames, passing them through a ResNeXt CNN architecture (Xie et al.,, 2016) that was pre-trained solely from the visual input from this child, and frozen except for a learnable linear projection head, to obtain a single embedding for each frame. CVCL’s language encoder is designed to take in transcribed child-directed utterances, passing each word through a single Embedding layer separately and then averaging across each word’s embedding in the utterance, again to obtain a single embedding for each utterance. Both encoders are jointly trained by passing embeddings into a contrastive loss, bringing closer matched embeddings from paired video frames and utterances, while separating mismatched embeddings from different video frames and utterances, similar to CLIP (Radford et al.,, 2021). This process of joint associative and representation learning produces a shared multimodal representation space, where neural similarity between words and referents determines the strength of their mapping.
We systematically explore three different architectural modifications to CVCL to examine their effects on grounded word learning. Each of these variants utilize the same underlying vision encoder architecture, but vary in the choice of language encoder and learning objective, each of which is illustrated in Figure  2.
3.1 Vision Encoder
For embedding static video frames, we replaced the ResNeXt-CNN in CVCL with a vision transformer (ViT-B/14) (Dosovitskiy et al.,, 2020), which is separately pre-trained only using visual data from the SAYCam dataset. This pre-training step is performed separately for each child’s visual data, using the self-supervised approach from DINO (Orhan and Lake,, 2024; Caron et al.,, 2021). From the set of frames associated with a given utterance (up to 16 sampled frames), we always sample one frame randomly to be paired with the corresponding utterance. During training, data augmentation is performed by performing a random crop to each video frame, followed by resizing the frame to $224\times 224$ pixels, and occassionally horizontally flipping the frame. During evaluation, frames are instead always resized to 256 pixels along the minor edge, followed by a $224\times 224$ centered crop.
We obtain the output of the vision transformer from the [CLS] token, a single embedding of size 768. This embedding is fed through a learnable linear projection head, followed by a layer norm and a dropout layer, such that the vision encoder returns a single embedding of size $D=512$. Similar to Vong et al., (2024), the bulk of the vision encoder is frozen during training, except for the learnable linear projection head.
3.2 Variant 1: Language encoder with minimal syntax (CVCL)
In this model variant, we examine whether providing additional word order information may be useful for grounded word learning compared to a bag-of-words approach, by incorporating positional information into the embedding. Because this is a relatively minor change from the original CVCL model, we still refer to this variant as CVCL throughout.
Language Encoder (Embedding): We used the same Embedding encoder from Vong et al., (2024), which extracts separate embeddings for each word in an utterance, and computes the average across all word embeddings to produce a single utterance embedding of size $D=512$. In this variant, we add a learned, absolute positional embedding to each token embedding before the averaging step. Prior to passing utterances into the language encoder, tokenization was performed using a word-level tokenizer built via spaCy.
Figure 3: Visualization of the contrastive objective. Given multiple pairs of image frames and utterances, we obtain embeddings via their respective encoders. The contrastive objective aims to maximize the cosine similarity of these matched embeddings (shown in green), while also minimizing the cosine similarity of mismatched embeddings (shown in red).
Learning Objective (Contrastive): Following a standard approach to training vision-language models (Radford et al.,, 2021; Jia et al.,, 2021), we employ an InfoNCE contrastive learning objective, which aims to bring together frame embeddings with their corresponding utterance embeddings in a mini-batch (that is, encouraging a larger dot product between the vectors), while separating all other mismatching frame and utterance embeddings (smaller dot product between the vectors). First, we obtain visual embeddings for each frame, $\mathbf{v_{i}}=f_{\theta}(\mathbf{x_{i}})$ as well as corresponding text embeddings, $\mathbf{u_{i}}=f_{\phi}(\mathbf{w_{i}}).$ Then, the contrastive objective aims to match each frame with its corresponding utterance:
$L_{\text{frame}}=-\frac{1}{N}\sum_{i}^{N}\log\frac{\exp(\mathbf{v_{i}}^{T}\mathbf{u_{i}}/\tau)}{\sum_{j=1}^{N}\exp(\mathbf{v_{i}}^{T}\mathbf{u_{j}}/\tau)},$ (1)
as well as matching each utterance with its corresponding frame:
$L_{\text{utterance}}=-\frac{1}{N}\sum_{i}^{N}\log\frac{\exp(\mathbf{u_{i}}^{T}\mathbf{v_{i}}/\tau)}{\sum_{j=1}^{N}\exp(\mathbf{u_{i}}^{T}\mathbf{v_{j}}/\tau)}.$ (2)
These two losses are then summed together and averaged:
$L_{\text{contrastive}}=\frac{1}{2}L_{\text{frame}}+\frac{1}{2}L_{\text{utterance}}.$ (3)
The contrastive objective is visualized in Figure 3.
3.3 Variant 2: Language encoder via Transformer (CVCL+T)
In this variant, we examine the effect using a more powerful language encoder. In the previous variant, the language encoder treats the importance of each word in an utterance equally, regardless of its context. To test this, we replace the Embedding-based language encoder with a 2-layer Transformer Decoder architecture instead, enabling it to create contextualized word representations by dynamically weighting contributions from surrounding words. We refer to this variant as CVCL+T (where +T indicates the use of the Transformer-based language encoder).
Language Encoder (Transformer): We used a GPT-2 transformer decoder as our language encoder (Radford et al.,, 2019). Since GPT-2 uses a causal self-attention mask, the embedding from the last hidden state contains a contextualized aggregate representation of all of the preceding words in the utterance, allowing us to treat this representation as an utterance encoding that is further refined via contrastive learning with the vision encoder. All GPT-2 transformer models use 2 layers, 8 attention heads, and shared weights for the embedding and unembedding layers. We use the embedding of the last hidden state (corresponding to the <eos> token) as the output embedding, which is also of size $D=512$.
Learning Objective (Contrastive): This variant utilizes the same contrastive learning objective as Variant 1.
3.4 Variant 3: Language encoder via Transformer and language modeling
In this variant, we examine the effect of exposing the model to richer linguistic signals via an auxiliary language modeling objective. The original contrastive objective purely aims to maximize visual-text alignment, but may be missing out on some of the other statistical regularities present in the structure of language. To test this, we augment the contrastive objective from CVCL+T with an auxiliary language modeling objective, training the model to predict each next word in an utterance. We refer to this variant as CVCL+T+LM throughout the paper, where the +LM indicates the addition of the language modeling objective to the base Transformer architecture.
Language Encoder (Transformer): This variant utilizes the same 2-layer Transformer architecture as as Variant 2 as its language encoder. For the variant, we also separately utilize the decoder’s logits for predicting the next word in a given utterance.
Learning Objective (Contrastive with Language Modeling): In addition to the contrastive learning objective, for this variant, we incorporate an additional language modeling loss, similar to (Zhuang et al.,, 2024). The equation is given by:
$L_{\text{joint}}=L_{\text{LM}}+\lambda_{c}L_{\text{contrastive}}$ (4)
We also scale the contrastive loss relative to the language modeling loss by $\lambda_{C}$, and through a hyperparameter sweep, we set it to be 0.3 across all experiments (Zhuang et al.,, 2024).
3.5 Training Details
Each model was trained up to 100 epochs using the AdamW optimizer, with a learning rate of 1e-4 or 1e-5 depending on the model type. The learning rate schedule involved 5000 steps of linear warmup, followed by cosine annealing until the end of training. Early stopping was performed using the validation loss. [^3] To better track learning, we filter each validation set with CLIP to only contain frame-utterance pairs with high CLIP similarity scores. See Appendix B for additional details. For the vision encoder, we followed the specifications in Vong et al., (2024); Orhan and Lake, (2024): we used the the augmentations as in the vision encoder pretraining (described above), we sampled a randomly associated frame with each utterance, and we froze the vision encoder during training (except for the projection head). For the language encoder, we use a maximum sequence length of 48. To train the model, we use a fixed temperature parameter in the contrastive loss with $\tau=0.07$, with varying batch sizes depending on the model configuration. Vision dropout, language dropout and weight decay were all set to 0.1. We performed hyperparameter search over batch size, learning rate, vision dropout, language dropout, weight decay, and $\lambda_{c}$. For each dataset and model configuration, we trained 3 separate models with different random seeds. A full comparison between the various the architecture and training details is shown in Table 3.
4 Results
One of the main evaluation methods for the original CVCL model was examining its zero-shot classification performance, specifically using the Labeled-S evaluation dataset, as described in Section 2.2. This evaluation dataset consists of 22 different categories, with 100 4-way classification trials per category. Previously, Vong et al., (2024) showed that an Embedding-based multimodal neural network (CVCL) trained on the subset of manually transcribed data from baby S achieved 61.6% classification accuracy on this dataset. In this section, we compare and contrast the performance of the original CVCL model against additional variants of CVCL trained using Whisper-derived transcriptions.
4.1 Finding 1: Comparable performance for models trained with automatic transcriptions vs. manual transcriptions
Figure 4: Labeled-S Classification Performance between CVCL trained with manual transcriptions (S-Manual-2022) vs. trained with Whisper transcriptions (S-Whisper-2022). Note that version of CVCL in the right plot utilizes a pre-trained vision transformer (ViT) instead of the CNN (ResNeXt), as well as a positional encoding scheme in the language encoder. Models were trained with three random seeds and results were averaged.
We first evaluated whether automatic transcriptions from Whisper would be comparable to manual transcriptions for training multimodal neural networks. These models were trained using a subset of Whisper transcribed utterances belonging to the same subset of videos that were previously manually transcribed and used to train CVCL, enabling a direct comparison of Whisper’s transcription quality. While initial analyses showed that the word frequency of the 22 categories from the Labeled-S evaluation showed a high correlation between the manually transcribed vs. Whisper transcriptions (Zhang et al.,, 2025), this did not provide any information about the quality of the Whisper-detected timestamps nor their temporal alignment with visual referents present in the dataset.
As can be seen in Figure 4, our results show that training CVCL on Whisper transcriptions provides comparable performance to CVCL, achieving a classification accuracy of 56.9%, compared to 61.6% from the original model. While there was a slight drop in performance compared to the original CVCL model, our results show that word-referent mappings are indeed learnable from Whisper-transcribed utterances, despite some of its limitations previously observed in transcription quality (Zhang et al.,, 2025). There are two additional reasons why the quality of Whisper transcriptions may have been sufficient to reach a similar level of performance as CVCL. First, rather than requiring perfect transcriptions for every word from every child-directed utterance, we only evaluate on a subset of the most prominent and commonly spoken nouns, for which there were a sufficient number of examples to learn from. Second, because of the relatively slow timescales at which visual information changes from the child’s perspective, large differences between the between relevant visual referents and detected utterance timings may not have been as important, something that was also observed with CVCL due to how some of its utterances were pre-processed during training.
4.2 Finding 2: Comparable performance for models trained on disjoint transcriptions
Figure 5: Labeled-S Classification Performance between CVCL trained on different combinations of Whisper transcriptions. The left-most bar (S-Whisper-2022, same as Figure 4), represents classification performance when trained on transcriptions the set of previously manually transcribed videos. The middle bar (S-Whisper-Disjoint) represents performance when trained only using transcriptions from new and non-overlapping videos from baby S. Finally, the right-most bar (S-Whisper) represents performance when trained using all of the combined Whisper-transcribed data from original and new videos. Models were trained with three random seeds and results were averaged.
Next, to examine whether our model genuinely learns generalizable word-referent mappings, we conducted a follow-up experiment where the frames for training versus evaluation are completely disjoint. This contrasts with the previous experiment, where training and evaluation frames were independently extracted, but from the same set of videos. We trained a distinct set of models using the S-Whisper-Disjoint split, which utilizes Whisper transcriptions derived from an independent and non-overlapping set of videos from baby S, distinct from the videos used in both the S-Whisper-2022 videos and the Labeled-S evaluation set. Since all of the training data in this split comes from entirely new videos from baby S that were not previously manually transcribed, this enables us to examine whether the learned word-referent mappings generalizes to the Labeled-S dataset. In this instance, the frames from Labeled-S are now considered as new, previously unseen visual instances of these categories that were not seen at all during training for the S-Whisper-Disjoint models, and can be thought of as a form of cross-dataset generalization. This kind of exploration was not previously possible due to the limited available set of manually transcribed videos that were used for both training and evaluation (although the construction of each respective dataset was independent).
The results are shown in Figure 5 (middle bar), and here we find that CVCL, when trained on non-overlapping data, also achieves a classification accuracy of 57.1%, indicating a relatively minimal drop in classification accuracy from generalizing within-distribution to generalizing out-of-distribution. Why do these models generalize well in this cross-dataset fashion? Despite being trained on a different set of videos, fundamentally, they are drawn from the same distribution of videos captured from a head-mounted camera in the same environment as the original manually transcribed videos for baby S, with similar objects and utterances across different videos. However, it is also the case that the contrastive objective forces multimodal representations to be robust and generalizable (Radford et al.,, 2019; Vong et al.,, 2024), guiding the models to learn the semantic relationship for different word-referent mappings, enabling broad generalization rather than overfitting or memorizing specific instances.
4.3 Finding 3: Classification performance increases with additional training data
Next, to determine whether increased training data could improve model performance, we trained models that used the maximum amount of training data available from baby S. These models were trained on the S-Whisper split, which was constructed using a combination of the Whisper transcriptions from S-Whisper-2022 and S-Whisper-Disjoint, resulting in approximately 3.5 times more data than used to train the original CVCL model (see Table 1 for additional details).
As shown in Figure 5 (right bar), we find that a CVCL model trained on this combined dataset split achieves an overall classification accuracy of 63.2%. While this outperforms CVCL, the gains in accuracy are relatively minimal given the large relative increase in training data. There are several possible explanations for the limited benefit to more training data. One reason similar to what was discussed in the previous section, the additional data from the newly transcribed videos in S-Whisper-Disjoint are being drawn from the same overall distribution of scenes and utterances, and therefore only providing a modest boost to the model’s generalization capabilities. An alternative explanation is that some categories from the Labeled-S evaluation set are sufficiently noisy or challenging, not just for CVCL or the variants we explore here, but for all models. For instance, Vong et al., (2024) evaluated CLIP ViT-L/14 on this dataset and found its zero-shot classification performance was only marginally higher (66.7%) despite being a strong vision-language foundation model, with CLIP also performing relatively poorly for a number of ambiguous categories like “floor” and “hand”.
4.4 Finding 4: Transformer-based language encoders achieve similar levels of classification performance
Next, to examine whether more sophisticated language encoders could improve grounded word learning, we replaced the Embedding-based language encoder with two variants involving a 2-layer Transformer. One of these variants used the same contrastive loss alone (CVCL+T), while the other variant combined this contrastive loss with an additional language modeling loss (CVCL+T+LM). These additional language encoder variants were chosen to examine whether other approaches for encoding utterances would be helpful for grounded word learning. The results in Figure 6 show that the Transformer-based language encoders show a similar or slightly worse performance compared to CVCL across the different splits, although no particular variant is consistently better or worse. One reason we may not have seen any additional advantages from stronger contextual-based encoding is the method of evaluation in our image classification tasks, which relies on only passing in the target category as a single token to the model during evaluation, rather than any additional words as context as seen during training.
4.5 Finding 5: Robust evidence of learnability from training models across different children’s datasets
Figure 6: Labeled-S Classification Performance between different combinations of Whisper transcriptions, across three different model configurations. The left-most bar in each group represents performance from Figure 5. The two additional bars in each group explored modifications to the language encoder, either replacing the Embedding encoder with a Transformer (T) or with a Transformer and an additional language modeling head (T+LM). Models were trained with three random seeds and results were averaged.
Finally, to test whether our findings generalize beyond a single child’s linguistic environment, we examined model performance across multiple children’s datasets. This addresses one of the key limitations from our previous analyses, which focused exclusively on learning from just one child’s data. While it has been sufficient to demonstrate learnability from relatively generic learning mechanisms, it fails to address how generalizable and robust these findings are. Here, we leverage Whisper’s ability to generate automatic transcriptions for all babies from the SAYCam dataset: S, A and Y. In this analysis, we train models on either S-Whisper, A-Whisper, or Y-Whisper.
In order to properly evaluate these models trained on other children’s data, we also required a separate set of evaluation trials that were drawn from each child’s videos respectively. This set of evaluations also consists of 4-way classification trials, but generated from a larger set of 41 categories, see Appendix C for additional dataset creation details. Notably, these evaluations were designed in a similar fashion to the S-Whisper-Disjoint split described above: the evaluation frames were generated from the “test” splits of each child’s data, such that they were independent from anything seen during training or validation. [^4] However, due to differences in dataset splitting, the the new set of evaluation frames derived from baby S are not completely independent from training on S-Whisper, but they are for the A-Whisper and Y-Whisper datasets. These evaluation datasets all contain the same set of 41 categories that were present in all three children’s visual and linguistic environments, enabling both a more comprehensive and comparable method of comparison for grounded word learning across children. Furthermore, we generate a different evaluation dataset from each child’s test frames and only evaluate models trained on that child’s data to their corresponding evaluation set, which we call Labeled-S-V2, Labeled-A and Labeled-Y respectively.
Results for these evaluations are shown in Figure 7. At the highest level, we find that all model configurations and dataset splits, when evaluated on their corresponding evaluation split, achieves a classification accuracy higher than random chance (25%). Overall, these results show that word-referent mappings are robustly learnable with the use of multimodal neural networks, not just in the case of one child’s dataset (Vong et al.,, 2024), but instead across multiple independent children’s datasets and across a variety of different model configurations. Looking more closely at the performance across dataset splits, we still see some qualitative differences in performance. Models trained on data from baby S and baby Y achieve higher levels of accuracy (at 51% and 49% respectively), while the best model configuration for baby A only achieves 43% accuracy. Similar to the previous set of results in comparing model variants on baby S’s dataset splits, here we also find that no specific model configuration achieves superior classification performance, but instead performance is generally in the same range for each split. Overall, these findings provide strong evidence that multimodal neural network based approaches can be used to acquire word-referent mappings, both for a larger set of categories than previously explored, and across models trained from different children’s datasets.
Figure 7: Classification accuracy across models trained from Whisper transcriptions of each child in SAYCam. Each model was evaluated on a new set containing 41 different categories, shared across the three evaluation datasets but using video frames derived from each child’s videos (see Figure 1 for some example evaluation frames). Models were trained with three random seeds and results were averaged.
5 Discussion
In this work, we explored the robustness of simulated word learning from developmentally plausible datasets using multimodal neural networks with minimal inductive biases. We used automatic speech transcription methods to transcribe the entire SAYCam dataset, increasing the number of available transcribed utterances by a factor of 7, and use these automatically transcribed utterances to generate a number of multimodal training and evaluation datasets. We tested 3 different model configurations varying the type of language encoder and loss functions, against 5 different dataset splits across 3 babies. Overall, our results showed that all model configurations, across all three datasets, demonstrated that word-referent mappings were learnable using multimodal neural networks. Our findings build upon the results from Vong et al., (2024) — although here we used almost twice as many visual concepts — highlighting that the kind and quantity of visual and linguistic information present in children’s environments is indeed sufficient to acquire and generalize many word-referent mappings, without the need for additional linguistic inductive biases. Across multiple dataset splits, we show that the learned word embeddings generalize to visual instances to entirely new videos not seen during the training process.
While the use of automatic transcriptions for multimodal modeling has previously been explored (Roy and Pentland,, 2002; Harwath et al.,, 2018), it has seen less uptake in the field of language acquisition, likely due to some of the difficulties in obtaining accurate and representative transcripts (although see Long et al., (2024) for another recent example). In this work, we find that training multimodal models from automatic transcripts enabled the training of separate models for each of the three children’s dataset from the SAYCam corpora, that were previously inaccessible due to the efforts required for manual speech transcription. Furthermore, by comparing models trained from the same subset of videos from baby S either from manual vs. automated transcriptions, we find only a minor gap in classification performance. However, one advantage in our particular case for examining word-referent mappings is that the visual input of children changes relatively slowly given the egocentric nature of these datasets. Thus, Whisper’s imperfect utterance timings, combined with the wide range from which we sample paired video frames, is apparently sufficient for learning.
While our results showed that the specific model configuration used for each dataset resulted in minor variations in classification performance, we found that no specific model configuration was dominant across all datasets. However, in the second set of evaluations, comparing performance across models trained across the three children with the same set of evaluation categories (although with different frames), we observed that performance on models trained on baby A were much worse compared to babies S or Y. One hypothesis we arrived at after watching a number of the videos is that there are far fewer instances of visual-linguistic alignment present in the training set. In particular, baby A’s recordings (and transcribed utterances) often involved both parents having a discussion that was unrelated to the captured visual input, providing limited instances for multi-modal alignment for the model. While prior research has shown that children can acquire new words from overhearing, they often still required the novel words to be relevant in the current visual context (Akhtar,, 2005; Gampe et al.,, 2012).
Overall, our results provide a roadmap for studying multimodal language acquisition in developing children. Compared to Vong et al., (2024), we explore datasets from multiple children and numerous model types and configurations, showing robustness to the acquisition of word-referent mappings across all of these configurations. Nevertheless, there still remain many other interesting research opportunities at the intersection of developmentally rich datasets and modern machine learning methods to further close the data-efficiency gap and understand how humans effectively learn language from so little. In particular, we hope future work will build on our learnability results to develop more psychologically and developmentally plausible cognitive models (Portelance and Jasbi,, 2024), incorporating more realistic constraints as well as mechanisms for learning from actions and social interaction.
Acknowledgments
We are grateful to the authors of the SAYCam article (Sullivan et al.,, 2022), and the volunteers who contributed to the data set. Thanks to Emin Orhan for training separate vision transformers from each child’s data, which were used as the vision backbone for our multimodal neural networks. Additional thanks to Qiwen Zhang, Hailey He, Manli Zhao, Luyang Shang and Michael Picheny for their guidance on using Whisper and WhisperX to help automatically transcribe the SAYCam dataset, and their efforts to manually verify the quality of the transcriptions for this work.
Appendix A Generating Evaluation Datasets
In this work, we generated three new evaluation datasets (Labeled-S-V2, Labeled-A and Labeled-Y) to evaluate our multimodal models, which was performed using a mixture of both automatic and manual filtering. To accomplish this, we created new labeled datasets derived from the corresponding video frames from (primarily) the test set videos for each child, [^5] The test set videos were unused during the training procedure, although for baby S there was some partial overlap for baby S creating three evaluation datasets consisting of 41 shared categories with 100 evaluation trials each, for a total of 4100 tirals.
These evaluation sets were created by first generating a set of target words or concepts to filter for. We generated this list by first considering the set of single word concepts from the MacArthur-Bates Communicative Development Inventory (MCDI), followed by filtering to (1) only consider words whose concreteness score was higher than 4.8 (using concreteness scores from Brysbaert et al., (2014)), (2) present in all three children’s vocabularies, and (3) whose part-of-speech tag was either a NOUN or VERB, producing a set of 156 potential target labeled concepts to scan for.
Next, we used the largest available CLIP model (ViT-L/14@336px) (Radford et al.,, 2021), and computed a similarity score (via cosine similarity) between every frame from each child’s set of test set frames and the set of 156 target concepts (using the text embedding of the target label directly), providing us with an automated method to detect potential labeled frames. Using a similarity threshold of 0.24 (based on qualitative observations), we selected the highest scoring label for any given video frame above this threshold, which provided us with a set of labeled frames for each child. Using this as a starting point, we then performed a second round of manual filtering, by examining every automatically labeled instance to determine whether it was valid or not for the corresponding category label. After this manual filtering step, we performed a final filtering step to only include categories for which there were 10 or more instances in each of the three children’s labeled datasets, leading to 41 categories in total.
Finally, we generated separate evaluation datasets for each child by sampling 100 trials per category with these corresponding labeled frames paired with three other randomly selected frames from other randomly selected categories from the same child, in a similar manner to the creation of the Labeled-S evaluation dataset, leading to three separate evaluation datasets consisting of 4100 trials each.
Appendix B Generating the CLIP-filtered Validation Sets
Due to the limited signal from the noisy frame-utterance pairs that were automatically extracted, we needed a stronger signal to determine whether models were learning during the training process. To facilitate this, we leveraged CLIP to separately create filtered versions of the validation set for each data split, which was performed once as a pre-processing step prior to model training, ensuring that only frame-utterance pairs whose similarity exceeded a threshold were used to compute the validation loss. We used CLIP ViT-B/32 (Radford et al.,, 2021), calculating the cosine similarity between a randomly selected frame and its associated utterance. We filtered the validation set to only include frame-utterance pairs whose similarity scores were greater than 0.24, [^6] This threshold was determined via manual inspection from a small number of similarity scores. thereby removing noisy or uninformative pairs for model validation during the training process. A comparison table for the size of the CLIP-filtered validation sets for each dataset split can be found in Table 2.
Appendix C Evaluation Dataset Categories
Labeled-S Categories (22 categories)
ball, basket, car, cat, chair, computer, crib, door, floor, foot, ground, hand, kitchen, paper, puzzle, road, room, sand, stairs, table, toy, window
Labeled-S-V2, Labeled-A and Labeled-Y Categories (41 categories)
arm, ball, banana, bed, bedroom, blanket, book, boots, bottle, bowl, box, bread, bucket, cereal, cow, cup, door, finger, foot, hair, horse, jeans, kitchen, knee, leg, monkey, napkin, oven, paper, pen, pillow, porch, shirt, shoe, shoulder, sock, sun, table, towel, toy, window
Appendix D Tables
Dataset Train Validation Test Total S-Whisper-2022 Number of videos 348 21 17 386 Number of utterances 39649 2348 2045 44042 Avg. utterance length 4.76 4.97 4.70 4.77 Number of extracted frames 363782 22509 19347 405638 Avg. frames per utterance 9.18 9.59 9.46 9.21 Total words 188610 11674 9621 209905 Vocabulary size - - - 2268 S-Whisper-Disjoint Number of videos 623 100 286 1009 Number of utterances 91953 10327 33715 135995 Avg. utterance length 4.57 4.84 4.74 4.64 Number of extracted frames 799615 98097 307541 1205253 Avg. frames per utterance 8.70 9.50 9.12 8.86 Total words 420609 50024 159881 630514 Vocabulary size - - - 3375 S-Whisper Number of videos 916 46 47 1009 Number of utterances 122571 6888 6536 135995 Avg. utterance length 4.64 4.59 4.60 4.64 Number of extracted frames 1088036 60539 56678 1205253 Avg. frames per utterance 8.88 8.79 8.67 8.86 Total words 568852 31606 30056 630514 Vocabulary size - - - 3856 A-Whisper Number of videos 299 38 44 381 Number of utterances 72810 9357 8810 90977 Avg. utterance length 4.65 5.10 4.65 4.69 Number of extracted frames 689794 90374 85755 865923 Avg. frames per utterance 9.47 9.66 9.73 9.52 Total words 338476 47698 40949 427123 Vocabulary size - - - 4199 Y-Whisper Number of videos 246 34 31 311 Number of utterances 71101 8688 8816 88605 Avg. utterance length 4.84 4.55 4.85 4.81 Number of extracted frames 670820 82276 85271 838367 Avg. frames per utterance 9.43 9.47 9.67 9.46 Total words 344221 39511 42747 426479 Vocabulary size - - - 4238
Table 1: Dataset Descriptives
Dataset Val CLIP-Filtered Val
S-Whisper-2022
Utterances 2,348 350
Total Frames 22,509 3,478
S-Whisper-Disjoint
Utterances 10,327 1,602
Total Frames 98,097 16,482
S-Whisper
Utterances 6,888 932
Total Frames 60,539 8,899
A-Whisper
Utterances 9,357 925
Total Frames 90,374 9,494
Y-Whisper
Utterances 8,688 870
Total Frames 82,276 9,142
Table 2: Comparison of Standard vs. CLIP-Filtered Validation Splits
Model CVCL Vong et al. (2024) CVCL CVCL (T) CVCL (T+LM) Architecture Vision encoder ResNeXt-50 ViT-B/16 ViT-B/16 ViT-B/16 Embedding size 512 512 512 512 Language encoder Embedding Embedding Transformer Decoder Transformer Decoder Number of attention heads - - 8 8 Number of layers - - 2 2 Positional encoding - Learned, absolute Learned, absolute Learned, absolute Max sequence length 25 48 48 48 Frame sampling method Sample single frame Sample single frame Sample single frame Sample single frame Pre-trained vision encoder True True True True Fine-tune vision encoder False False False False Image data augmentation True True True True Loss Configuration Loss Contrastive Contrastive Contrastive Contrastive and Language Modeling Fixed temperature True True True True Initial temperature 0.07 0.07 0.07 0.07 $\lambda_{\text{Contrastive}}$ 0.3 Training Configuration Train batch size 8 16 64 64 Validation batch size 8 8 8 8 CLIP-filtered Validation False True True True Optimizer AdamW AdamW AdamW AdamW Learning rate 1e-4 1e-4 1e-5 1e-5 LR Scheduler ReduceLROnPlateau Linear Warmup with Cosine Annealing Linear Warmup with Cosine Annealing Linear Warmup with Cosine Annealing Warmup steps - 5000 5000 5000 Weight decay 0.1 0.1 0.1 0.1 Vision dropout - 0.1 0.1 0.1 Vision layer norm - True True True Language dropout - 0.1 0.1 0.1 Num epochs 400 100 100 100 Early Stopping Validation loss Validation loss Validation loss Validation loss
Table 3: Comparison of model configurations